If x-y=6 and 3x(squared)+3xy-6y(squared)=60, then 9x+18y=?
Answer by jaymore(5)

3xsquared+3xy-6ysquared=60...Factorise
3x(x+2y)-3y(x+2y)=60
(3x-3y)(x+2y)=60....1
3x-3y can be written as 3(x-y)...2
Substitute x-y=6 in 2 above
Now 3x-3y=18
go back to 1 above and plug in 18

18(x+2y)=60....Divide both sides by 18
x+2y=60/18........3
9x+18y can be written as 9(x+2y)
9(x+2y)=9times60/18=30